### v2

# EvoAgentX: An Automated Framework for Evolving Agentic Workflows

Abstract
Multi-agent systems (MAS) have emerged as a powerful paradigm for orchestrating large language models (LLMs) and specialized tools to collaboratively address complex tasks. However, existing MAS frameworks often require manual workflow configuration and lack native support for dynamic evolution and performance optimization. In addition, many MAS optimization algorithms are not integrated into a unified framework. In this paper, we present EvoAgentX, an open-source platform that automates the generation, execution, and evolutionary optimization of multi-agent workflows. EvoAgentX employs a modular architecture consisting of five core layers: the basic components, agent, workflow, evolving, and evaluation layers. Specifically, within the evolving layer, EvoAgentX integrates three MAS optimization algorithms, TextGrad, AFlow, and MIPRO, to iteratively refine agent prompts, tool configurations, and workflow topologies. We evaluate EvoAgentX on HotPotQA, MBPP, and MATH for multi-hop reasoning, code generation, and mathematical problem solving, respectively, and further assess it on real-world tasks using GAIA. Experimental results show that EvoAgentX consistently achieves significant performance improvements, including a 7.44% increase in HotPotQA F1, a 10.00% improvement in MBPP pass@1, a 10.00% gain in MATH solve accuracy, and an overall accuracy improvement of up to 20.00% on GAIA. The source code is available at: https://github.com/EvoAgentX/EvoAgentX.
EvoAgentX: An Automated Framework for Evolving Agentic Workflows
Yingxu Wang1 Siwei Liu2 Jinyuan Fang3 Zaiqiao Meng3††thanks: Corresponding author. 1Mohamed bin Zayed University of Artificial Intelligence 2University of Aberdeen 3University of Glasgow yingxv.wang@gmail.com, siwei.liu@abdn.ac.uk j.fang.2@research.gla.ac.uk, zaiqiao.meng@glasgow.ac.uk
1 Introduction
Multi-agent systems (MAS) are emerging as a powerful paradigm for orchestrating large language models (LLMs) and specialized tools to solve complex tasks collaboratively Hong et al. (2023); Gao et al. (2024); Fang et al. (2025). By coordinating multiple agents with distinct capabilities, such as planning, reasoning, or code generation, MAS decompose intricate problems into controllable subtasks and assign them to agents capable of solving them Yuan et al. (2024); Zhang et al. (2025a). This flexible and modular architecture makes MAS well-suited for addressing complex real-world problems. As a result, MAS have been widely deployed in applications such as multi-hop question answering Hong et al. (2023), software engineering automation Li et al. (2023), code generation Liu et al. (2025), mathematical problem solving Gao et al. (2024), and dialogue systems Shi et al. (2024).
Despite these promising developments, constructing a multi-agent system typically requires manual efforts in defining the roles of agents, specifying task decomposition strategies, designing agent interactions, and configuring execution workflows Tang et al. (2024); Xiao et al. (2024). Frameworks like CrewAI 111https://www.crewai.com/, CAMEL AI Li et al. (2023), and LangGraph 222https://www.langchain.com/langgraph have provided general-purpose frameworks for building such multi-agent systems, but they primarily rely on hand-crafted configurations or rule-based orchestration. This reliance limits scalability and usability, especially when adapting workflows to new tasks or domains. As the complexity of multi-agent systems grows, there is an urgent need for automating workflow construction to reduce manual efforts and facilitate the rapid development of agent-based applications.
Another challenge lies in enabling MAS to evolve and optimize themselves dynamically rather than relying on static, predefined configurations Zhang et al. (2025a). Real-world tasks, such as multi-hop question answering or software debugging, often involve changing inputs, intermediate outcomes, or increasing task complexities Xu et al. (2024). These characteristics make it essential for multi-agent systems to support workflow evolution and optimization to continuously adapt strategies to meet task requirements and respond to varying conditions. By adjusting workflow topologies, modifying prompt templates, and choosing the most suitable tools, MAS can iteratively refine their workflows for improved problem-solving effectiveness Zhou et al. (2024). However, most existing MAS frameworks lack mechanisms for dynamic workflow evolution or optimization. Although some multi-agent optimization methods like DSPy Khattab et al. (2023) and TextGrad Yuksekgonul et al. (2024) provide approaches for prompt refinement or agent orchestration, they remain fragmented and are not integrated into unified platforms, making it difficult for practitioners to apply and compare them consistently and effectively across diverse tasks.
To this end, we propose EvoAgentX, an open-source platform designed for the automated generation and evolutionary optimization of multi-agent workflows. As illustrated in Figure 1, EvoAgentX employs a modular architecture consisting of five core layers: basic components, agent, workflow, evolving, and evaluation. The central feature of EvoAgentX is the evolving layer, which seamlessly integrates three state-of-the-art optimization algorithms, TextGrad Yuksekgonul et al. (2024), AFlow Zhang et al. (2024b), and MIPRO Opsahl-Ong et al. (2024), to iteratively refine agent prompts, tool configurations, and workflow topologies. We evaluate EvoAgentX comprehensively across diverse benchmarks, including HotPotQA for multi-hop reasoning Yang et al. (2018), MBPP for code generation Austin et al. (2021a), MATH for mathematical problem-solving Hendrycks et al. (2021), and the GAIA benchmark for real-world multi-agent tasks Mialon et al. (2023). Experimental results demonstrate that EvoAgentX achieves substantial performance improvements through dynamic workflow evolution, including improvements of 7.44% in HotPotQA F1, 10.00% in MBPP pass@1, 10.00% in MATH solve accuracy, and up to 20.00% overall accuracy on GAIA.
The contributions of our demo are as follows:
We present EvoAgentX, an open-source platform that enables easy customization and automatic generation of multi-agent workflows from high-level goal descriptions, reducing manual efforts in system design.
EvoAgentX integrates three optimization algorithms, TextGrad, AFlow, and MIPRO, to enable evolutionary workflow optimization.
EvoAgentX provides built-in benchmarks and standardized evaluation metrics, supporting dynamic workflow evolution that consistently improves performance on datasets such as HotPotQA, MBPP, and MATH, as well as on real-world benchmarks like GAIA.
2 Related Work
2.1 Multi-Agent Systems
Recent research on multi-agent systems (MAS) has explored how multiple language model agents can collaborate to solve complex tasks by distributing subtasks across specialized components Wang et al. (2024); Kapoor et al. (2024). Frameworks such as CAMEL AI, CrewAI, and LangGraph provide general-purpose infrastructures that allow users to define agents with distinct roles, specify communication protocols, and design execution logic and interaction patterns. These systems have enabled significant progress in areas such as multi-hop reasoning, dialogue simulation, tool-augmented reasoning, and software engineering automation Ma et al. (2024). However, most existing MAS frameworks rely on hand-crafted workflows or rule-based orchestration, requiring users to manually predefine agent hierarchies, interactions, and execution strategies for each task Gao et al. (2024); Islam et al. (2024). This places a substantial burden on users and limits the scalability and adaptability of such systems, particularly when workflows need to be adjusted for new tasks or changing conditions. To address this limitation, EvoAgentX introduces automatic workflow generation from high-level task descriptions, eliminating the need for manual workflow design and significantly reducing human effort in building multi-agent systems.
2.2 Multi-Agent Optimization
Recent work on multi-agent optimization aims to enhance task performance by refining agent prompts, execution strategies, and communication structures Agarwal et al. (2024); Zhou et al. (2025). Early studies focused on prompt optimization, with methods such as DSPy Khattab et al. (2023) and TextGrad Yuksekgonul et al. (2024) demonstrating that prompt-level refinement improves orchestration efficiency. In parallel, frameworks like DyLAN Liu et al. (2023) and Captain Agent Song et al. (2024) explored dynamic topology adjustments through reactive modifications to agent sets or communication links. A significant advancement was achieved when multi-agent topology was formalized as a search and optimization problem. Approaches such as AutoFlow Li et al. (2024) and GPTSwarm Zhuge et al. (2024) defined topological search spaces using natural language programs or computational graphs and applied reinforcement learning to optimize agent connections. Building on this, ScoreFlow Wang et al. (2025) and G-Designer Zhang et al. (2024a) introduced preference learning and deep generative models to produce task-adaptive communication structures. Recent work has further unified prompt and topology optimization, as in ADAS Hu et al. (2024), FlowReasoner Gao et al. (2025), and MaAS Zhang et al. (2025b), which jointly optimize agent configurations, prompts, and execution graphs using meta-search, reinforcement learning, or probabilistic supernets. Despite these advances, most existing frameworks rely on fragmented toolchains or require manual setup, making them difficult to apply consistently across diverse tasks. EvoAgentX addresses this gap by providing an integrated platform that automates multi-agent workflow generation and unifies optimization techniques in a single end-to-end system.
3 System Design
EvoAgentX is an open-source framework designed to automate the generation, execution, evaluation, and evolutionary optimization of agentic workflows. It supports automatic multi-agent workflow generation from high-level task descriptions, seamless integration of evolutionary optimization algorithms such as TextGrad Yuksekgonul et al. (2024), AFlow Zhang et al. (2024b), and MIPRO Opsahl-Ong et al. (2024), and built-in benchmarks with standardized evaluation metrics for systematic performance assessment. To enable these capabilities, as shown in Figure 1, EvoAgentX adopts a modular architecture comprising five core layers: the basic components, agent, workflow, evolving, and evaluation layers.
3.1 Basic Component Layer
The basic components layer forms the foundation of EvoAgentX, providing essential services that ensure the system’s stability, scalability, and extensibility. It simplifies infrastructure management, supporting high-level agent design and workflow construction. Core modules include configuration management, logging, file handling, and storage. The configuration manager validates system parameters from structured files (e.g., YAML or JSON), while the logging module tracks system events and performance metrics. File handling components manage workflow states and agent checkpoints, ensuring experiment reproducibility. The storage manager supports both persistent and temporary storage, including caching and checkpointing.
To further enhance the system’s flexibility and adaptability, EvoAgentX integrates with a variety of LLMs through frameworks such as OpenRouter 333https://openrouter.ai/ and LiteLLM 444https://www.litellm.ai/ in the basic component layer, enabling seamless integration of LLMs from diverse sources.
3.2 Agent Layer
The agent layer serves as the core functional unit of EvoAgentX, enabling the construction of modular, intelligent entities that integrate reasoning, memory, and action execution capabilities in a seamless and flexible manner. Each agent is designed as a composition of a LLM, action modules, and memory components, together supporting flexible, context-aware decision-making and tools.
At the center of the agent architecture, the LLM is responsible for high-level reasoning, response generation, and context interpretation. It is specified through direct instantiation or configuration files and serves as the foundation for all agent operations. Actions define the operational logic of agents. Each action encapsulates a specific task (e.g., summarization, retrieval, API invocation) and consists of a prompt template, input-output format specifications, and optional tool integrations. Formally, an agent is represented as:
where denotes the memory module, and denotes the set of action components.
An example in A.1 illustrates how to create an agent within EvoAgentX.
3.3 Workflow Layer
The workflow layer is another core component of EvoAgentX, supporting the construction, orchestration, and execution of multi-agent workflows in a structured and flexible manner. It provides a formal representation for capturing task dependencies, execution flows, and communication between agents. Each workflow is modeled as a directed graph:
where denotes the set of nodes (tasks) and represents directed edges encoding dependencies and data flow between tasks. Each node corresponds to a WorkFlowNode, defining a specific task, its inputs, outputs, associated agents, and execution status (PENDING, RUNNING, COMPLETED, or FAILED). Nodes can encapsulate either a set of agents, allowing dynamic selection of optimal actions during execution, or an ActionGraph specifying an explicit sequence of operations. Edges capture task dependencies, execution order, and optional priority weights for scheduling.
The workflow layer supports both general-purpose workflows (WorkFlowGraph) and streamlined linear workflows (SequentialWorkFlowGraph). The former provides a flexible framework for explicitly defining complex task graphs, including custom nodes, edges, conditional branches, and parallel execution patterns. It allows users to specify detailed task dependencies and exercise fine-grained control over data flow and execution logic. In contrast, the latter is designed for simplicity, automatically inferring graph connections based on task input-output dependencies and generating nodes, agents, and edges without the need for manual graph specification. This dual design facilitates rapid prototyping while preserving the expressiveness needed to model complex structures.
An example in A.2 illustrates how to create a workflow within EvoAgentX.
3.4 Evolving Layer
The evolving layer of EvoAgentX consists of three core components: agent optimizer, workflow optimizer, and memory optimizer. These optimizers provide a unified mechanism for iteratively refining agent configurations, workflow topologies, and memory management strategies. This architecture enables the system to dynamically adapt to changing task requirements, optimize multi-agent coordination, and improve overall performance.
(1) The agent optimizer aims to refine agent prompt templates, tool configurations, and action strategies to enhance each agent’s performance across diverse tasks. Formally, for an agent parameterized by its prompt and configuration , the optimizer seeks to compute
where denotes the agent-level optimization operator that updates prompts and configurations based on evaluation feedback, and denotes evaluation feedback. The TextGrad Yuksekgonul et al. (2024) and MIPRO Opsahl-Ong et al. (2024) optimizers are employed for agent optimization, jointly applying gradient-based prompt tuning, in-context learning, and preference-guided refinement to iteratively align prompts, tool configurations, and agent outputs with task-specific objectives based on the performance signal.
(2) The workflow optimizer focuses on improving task decomposition and execution flow by adjusting the structure of the workflow graph . Formally, its objective is to compute
where denotes the workflow-level optimization operator that updates the graph structure based on evaluation feedback, and denotes evaluation feedback. The SEW Liu et al. (2025) and AFlow Zhang et al. (2024b) optimizers are employed for workflow optimization, iteratively restructuring workflow graphs by reordering nodes, modifying dependencies, and exploring alternative execution strategies guided by the task performance signal and convergence criteria.
(3) The memory optimizer remains under active development. Its objective is to provide structured, persistent memory modules that enable selective retention, dynamic pruning, and priority-based retrieval Zeng et al. (2024). Formally, it aims to compute
where denotes the memory-level optimization operator that updates the agent’s memory module based on evaluation feedback, and denotes evaluation feedback.
An example in A.3 demonstrates how to use the optimizer within EvoAgentX.
| Task | Dataset Name | # Train | # Dev | # Test |
| QA | NQ | 79,168 | 8,757 | 3,610 |
| Multi-Hop QA | HotPotQA | 90,447 | 7,405 | / |
| Math | GSM8K | 7,473 | / | 1,319 |
| Math | MATH | 7,500 | / | 5,000 |
| Code Generation | HumanEval | / | / | 164 |
| Code Generation | MBPP | / | / | 427 |
| Code Generation | LiveCodeBench (v1v5) | / | / | 400880 |
| Code Execution | LiveCodeBench | / | / | 479 |
| Test Output Prediction | LiveCodeBench | / | / | 442 |
3.5 Evaluation Layer
The evaluation layer of EvoAgentX provides a modular and extensible framework for systematically assessing workflow performance across tasks and benchmarks. It integrates two complementary components: (1) the task-specific evaluator and the LLM-based evaluator. The task-specific evaluator computes domain-relevant metrics by comparing workflow outputs against ground truth labels, supporting validation on datasets such as HotPotQA, MBPP, and MATH. More details about the benchmarks provided in EvoAgentX are shown in Table 1; (2) the LLM-based evaluator leverages the reasoning and generation capabilities of large language models to deliver flexible and context-aware evaluations. It supports qualitative assessments, consistency checking, and dynamic criteria that are not easily captured by static metrics.
Formally, given a workflow and dataset , the evaluation process is defined as
where maps workflow executions to aggregated performance metrics . The design supports both WorkFlowGraph and ActionGraph structures, allowing evaluation at various abstraction levels.
4 Experiments
We evaluate EvoAgentX across three tasks: (1) Evolution Algorithms, which optimize agent configurations and workflow topologies to improve performance; (2) Applications, where EvoAgentX is applied to enhance multi-agent systems on real-world benchmarks; and (3) Case Study, demonstrating EvoAgentX’s capability to optimize workflows and enhance agent performance through practical examples.
4.1 Evolution Algorithms
Experimental Settings. We evaluate EvoAgentX’s evolutionary optimization capabilities by applying three integrated algorithmsm, TextGrad, MIPRO, and AFlow, to iteratively refine agent prompts, tool configurations, and workflow topologies, using metrics such as F1 score for multi-hop reasoning, pass@1 accuracy for code generation, and solve rate for mathematical problem solving.
Datasets. We assess EvoAgentX on HotPotQA for multi-hop question answering requiring compositional reasoning Yang et al. (2018), MBPP for Python code generation from natural language descriptionsAustin et al. (2021b), and MATH for solving high-school level mathematical problems Hendrycks et al. (2021).
| Method | HotPotQA | MBPP | MATH |
| (F1%) | (Pass@1 %) | (Solve %) | |
| Original | 63.58 | 69.00 | 66.00 |
| TextGrad | 71.02 | 71.00 | 76.00 |
| AFlow | 65.09 | 79.00 | 71.00 |
| MIPRO | 69.16 | 68.00 | 72.30 |
Results Analysis. As shown in Table 2, EvoAgentX’s optimization algorithms consistently enhance performance across all benchmarks. Specifically, TextGrad substantially improves multi-hop reasoning, increasing the HotPotQA F1 score from 63.58% to 71.02%. AFlow significantly boosts code generation accuracy, raising MBPP pass@1 from 69.00% to 79.00%. Similarly, TextGrad strengthens mathematical reasoning, improving the MATH solve rate from 66.00% to 76.00%. These results demonstrate the EvoAgentX can effectively refine multi-agent workflow topologies to align with task-specific objectives across domains.
4.2 Applications
Experimental Settings. We apply EvoAgentX to optimize existing multi-agent systems on a real-world benchmark. Specifically, We select two representative open-source frameworks from the GAIA leaderboard Mialon et al. (2023), Open Deep Research 555https://github.com/langchain-ai/open_deep_research and OWL 666https://github.com/camel-ai/owl, and refine their agent prompts and workflow configurations using EvoAgentX, evaluating performance based on accuracy, which measures the correctness of generated answers against the ground truth.
Results Analysis. As shown in Figure 2, EvoAgentX significantly improves the performance of both Open Deep Research and OWL across all evaluation levels on the GAIA benchmark. For Open Deep Research, the overall accuracy increases by 18.41%, with notable improvements of 20.00% at Level 1, 8.71% at Level 2, and 7.69% at Level 3. Similarly, OWL achieves an overall accuracy improvement of 20.00%, driven by gains of 28.57% at Level 1, 10.00% at Level 2, and a remarkable 100.00% at Level 3. These results demonstrate the effectiveness of EvoAgentX in enhancing real-world multi-agent systems through automated prompt and topology optimization.
4.3 Case Study
We present a case study to illustrate the practical application of EvoAgentX in refining agent prompts and workflow configurations within existing multi-agent systems, including examples from AFlow, TextGrad, and MIPRO. More detailed analysis and results are presented in Appendix A.4.
5 Conclusion
We present EvoAgentX, an open-source platform that automates the generation, execution, evaluation and optimization of multi-agent workflows. It addresses key limitations of existing frameworks by eliminating the need for manual workflow design and providing support for dynamic, task-specific optimization. By integrating multiple optimization algorithms, including TextGrad, AFlow, and MIPRO, EvoAgentX enables the iterative refinement of agent prompts, tool configurations, and workflow topologies with minimal manual intervention. Experiments on diverse benchmarks, such as HotPotQA, MBPP, MATH and GAIA, demonstrate that EvoAgentX consistently achieves substantial performance improvements in multi-hop reasoning, code generation, mathematical problem solving, and real-world multi-agent applications. In the future, we will extend EvoAgentX with more optimization algorithms, richer tool integration, and long-term memory to further enhance agent adaptability and contextual awareness. We also plan to explore advanced evolution strategies Fang et al. (2025), including MASS Zhou et al. (2025), AlphaEvolve Novikov et al. (2025), and Darwin Gödel Machine Zhang et al. (2025c), to advance the strategies of multi-agent optimization.
Appendix A Appendix
A.1 Creating a Simple Agent in EvoAgentX
In EvoAgentX, a simple agent can be created using the CustomizeAgent class, which enables the quick configuration of agents with a specific prompt. To create such an agent, the first step is to configure the large language model (LLM) that will be used by the agent. This is done by defining the LLM settings, such as the model type and API key, using the OpenAILLMConfig class. After configuring the LLM, the agent is instantiated by specifying its name, description, and prompt, which defines the task the agent will perform.
The following code demonstrates the process of creating and using a simple agent in EvoAgentX, where the agent is tasked with printing "hello world." The agent is then executed, and the result is retrieved as a message object, with the content of the response extracted and displayed.
This approach showcases how a basic agent can be constructed and executed with minimal configuration in EvoAgentX.
A.2 Creating a Simple Workflow in EvoAgentX
In EvoAgentX, workflows enable multiple agents to collaborate sequentially on tasks. A basic sequential workflow involves defining tasks, each with a name, description, input-output specifications, and a prompt. To create such a workflow, the SequentialWorkFlowGraph class is used to define and link tasks. An agent manager oversees the agents responsible for executing each task. The workflow is executed by providing the necessary inputs, and the results are collected as outputs.
The following code illustrates the process of creating and using a simple workflow in EvoAgentX, where the workflow includes two tasks: "Planning," in which the agent creates a detailed implementation plan for a given problem, and "Coding," in which the agent implements the solution based on the plan. The workflow is then executed with a specified problem, and the results are retrieved as outputs, with the content of each task’s result extracted and displayed sequentially.
This example demonstrates how a simple sequential workflow can be created and executed in EvoAgentX, where agents collaborate in a defined sequence to accomplish a task.
A.3 A Simple Example about using Optimizer within EvoAgentX
In EvoAgentX, the AFlow optimizer optimizes multi-agent workflows for tasks like code generation. To use the optimizer, the first step is configuring the LLMs for both optimization and execution, one for optimizing the workflow (e.g., Claude 3.5 Sonnet) and one for task execution (e.g., GPT-4o-mini). The task configuration specifies operators (e.g., Custom, CustomCodeGenerate, ScEnsemble), and the workflow is created using the SequentialWorkFlowGraph class. The optimizer is initialized with the paths to the workflow, LLM configurations, and optimization parameters.
The following code demonstrates the process of creating and using an AFlow Zhang et al. (2024b) optimizer in EvoAgentX, where the optimizer refines a multi-agent workflow for code generation. The optimizer is configured with specific settings, including the selection of LLMs for both optimization and execution. The process begins by configuring the optimizer_llm for optimizing the workflow and the executor_llm for executing the tasks. The optimizer is run with a benchmark (e.g., the HumanEval benchmark), and the results are retrieved as outputs, with each optimization step and its corresponding result displayed sequentially.
This illustrates how EvoAgentX dynamically optimizes workflows to improve performance through multiple evaluation rounds.
A.4 Case study
A.4.1 AFlow for Workflow Optimization
In this task, we aim to enhance the efficiency and effectiveness of multi-agent workflows for mathematical problem-solving by optimizing workflows using AFlow, as shown in the Example 1 and Example 2.
The initial workflow used a basic agent configuration to solve mathematical problems with a simple prompt. After optimization with AFlow within EvoAgentX, the workflow was significantly enhanced by incorporating detailed problem analysis, Python code generation, and solution refinement through an ensemble approach. The workflow now involves multiple agents for problem analysis, code generation, and solution refinement, improving both accuracy and clarity. This optimization results in a more comprehensive and precise problem-solving process.
A.4.2 TextGrad for Prompt Optimization
In this task, we aim to improve the efficiency and effectiveness of multi-agent workflows for mathematical problem-solving by optimizing prompts using TextGrad, as demonstrated in Example 5 and Example 6.
The initial prompt employed a minimalistic instruction, directing the agent to solve mathematical problems by simply providing the final answer in a boxed format. Following optimization with TextGrad within EvoAgentX, the prompt was substantially refined to support a structured, step-by-step reasoning process. The enhanced prompt guides the agent to assess problem complexity, apply appropriate mathematical principles, and generate executable Python code. It further emphasizes logical coherence, justification of each solution step, and the use of explanatory transitions to enhance interpretability. These improvements lead to increased solution accuracy and significantly improve the clarity and transparency of the reasoning process.
A.4.3 MIPRO for Prompt Optimization
In this task, we aim to improve the efficiency and effectiveness of multi-agent workflows for mathematical problem-solving by optimizing prompts using MIPRO, as demonstrated in Example 5 and Example 6.
The initial prompt was designed to provide a mathematical solution in a simple boxed format, without elaborating on the solution steps or offering detailed explanations. Following optimization with MIPRO within EvoAgentX, the prompt was significantly improved to guide the agent through a comprehensive, step-by-step solution process. The enhanced prompt now incorporates intermediate steps, clear explanations of relevant mathematical concepts, and a thorough breakdown of the problem-solving approach. Each step is carefully articulated, ensuring a deeper understanding of the solution. This optimization not only improves the accuracy of the solution but also enhances the clarity and transparency of the reasoning process.